OpenStudy (anonymous):

What is the least integer n such that 1/(2^n) < 0.001 Answer options: 10 11 500 501 There is no such least integer On the GRE you can only use a very basic calculator that has no "power" button. How can I solve this without a fancy calculator?

OpenStudy (anonymous):

Out of those answers, the first one, 10, would be it. Without a fancy calculator, just do the math in your head. 2^n is simply doubling numbers. 2^0 = 1 2^1= 2 so just continue on... 2,4,8,16,32,64,128,256,512,1024 So it's 1/1024, which 1/1000 would be 0.001, so 1/1024 should be slight less. Perfect!

OpenStudy (anonymous):

\[ \frac 1{2^n} < \frac 1{1000}\\ 2^n > 10^3\\ n \ln(2) > 3 \ln(10)\\ n >\frac{ 3 \ln(10)}{\ln(2)}=9.96578\\ n=10 \]